OpenStudy (anonymous):

13. A ball is thrown upwards with a speed of 39.2 m/s. Calculate (a) the maximum height it reaches, and (b) the time taken in reaching the maximum height.

OpenStudy (anonymous):

Answers please tel

OpenStudy (abmon98):

U=39.2 a=-10m/s^2 as accelertion is opposite to gravity at maximum height v=0 Apply V^2-U^2=2as -(39.2)^2=2*-10*s -20s=-1536.64 s=76.832 meter v-u=at -39.2=-10t t=3.92 seconds